OpenStudy (anonymous):

Determine if the series is convergent Sum[n/e^n], n,1,inf.

OpenStudy (anonymous):

I would appreciate if you use the equation editor (latex) \[\large \sum_{1}^{\infty} \frac{n}{e^n}\] Is this your question ??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

If i used the ratio test, i think I get "1"

OpenStudy (anonymous):

You should not get 1 from the ratio test. I got 1/e from it.

OpenStudy (anonymous):

@bmp , could you show how to do this as I am new to this. Or just guide me how to do this?

OpenStudy (anonymous):

From the ratio test we have: \[ \lim_{n \rightarrow \infty} |\frac{n+1}{e^{n+1}} \frac{e^n}{n}| \implies | \frac{1}{e} \lim_{n \rightarrow \infty} \frac{n+1}{n} \]

OpenStudy (anonymous):

But the limit is equal to 1, so we are left with 1/e. Because the limit is exists and is < 1, we conclude that the series converges.

OpenStudy (anonymous):

Typo: the limit exists*

OpenStudy (anonymous):

@bmp Thanks a lot :D

OpenStudy (anonymous):

For some reason, I decided to make "e" magically disappear

OpenStudy (anonymous):

@shivam_bhalla No problem :-) I wasn't too careful with my typing tho. If something remained unclear, ask away.

OpenStudy (anonymous):

Thannks!!

OpenStudy (anonymous):

@bmp Just one thing. How did you get n+1/e^(n+1) in your answer. Is their any technique ??

OpenStudy (anonymous):

it's the ratio test rules

OpenStudy (anonymous):

It's the ratio test: you do \(\LARGE \lim_{n \rightarrow \infty} \frac{a_{n+1}}{a_n} \)

OpenStudy (anonymous):

It's the modulus of the division. If the limit exists and is < 1, the series converges. If the limit is > 1 or does not exist, the series diverges. If the limit is equal to 1, it's inconclusive.

OpenStudy (anonymous):

u forgot something bmp, it's the absolute of that

OpenStudy (anonymous):

Ok I got that, But just one question. If you are taking modulus then how can you arrive at a negative value??

OpenStudy (anonymous):

I mentioned on the second post. I don't recall the Latex for || by memory, haha.

OpenStudy (anonymous):

You can't. < 1 means that it will be between 0 and 1 or [0,1)

OpenStudy (anonymous):

Loks like I have to clear my glasses. :P I thought of it as less than 0 :P Thanks a lot @iHelp and @bmp

OpenStudy (anonymous):

*looks

OpenStudy (anonymous):

np, and bmp sorry- i missed the post

OpenStudy (anonymous):

That's fine. Had I forgot it, you would be perfectly correct and it would help shivam_bhalla. Anyway, glad to be helpful :-)